# Bootstrapping Language Models with DPO Implicit Rewards

### v1

@@ Abstract
Human alignment in large language models (LLMs) is an active area of research. A recent groundbreaking work, direct preference optimization (DPO), has greatly simplified the process from past work in reinforcement learning from human feedback (RLHF) by bypassing the reward learning stage in RLHF. DPO, after training, provides an implicit reward model. In this work, we make a novel observation that this implicit reward model can by itself be used in a bootstrapping fashion to further align the LLM. Our approach is to use the rewards from a current LLM model to construct a preference dataset, which is then used in subsequent DPO rounds. We incorporate refinements that debias the length of the responses and improve the quality of the preference dataset to further improve our approach. Our approach, named self-alignment with DPO ImpliCit rEwards (DICE), shows great improvements in alignment and achieves superior performance than Gemini Pro on AlpacaEval 2, reaching $27.55\%$ length-controlled win rate against GPT-4 Turbo, but with only 8B parameters and no external feedback. Our code is available at https://github.com/sail-sg/dice.
@@ 1 Introduction
Direct preference optimization (DPO) [17] presents a compelling alternative to reinforcement learning from human feedback (RLHF) in large language models (LLMs). By circumventing the complexity of learning a reward model from given human preferences, DPO is simpler to implement and train compared to the RLHF approaches. Importantly, DPO, once trained, implicitly specifies a reward model. Mathematically, the reward for a response $\mathbf{y}$ to the prompt $\mathbf{x}$ can be expressed in terms of the optimal policy $\pi^{\star}$ and the reference policy $\pi_{\textrm{ref}}$:
$\hat{r}(\mathbf{x},\mathbf{y})=\beta\log\frac{\pi^{\star}(\mathbf{y}|\mathbf{x})}{\pi_{\textrm{ref}}(\mathbf{y}|\mathbf{x})}+\beta\log Z(\mathbf{x})\textrm{,}$
@@
for parameter $\beta$ and normalizing constant $Z$. Further, an implicit reward $r(\mathbf{x},\mathbf{y})=\beta[\log\pi^{\star}(\mathbf{y}|\mathbf{x})-$ $\log\pi_{\textrm{ref}}(\mathbf{y}|\mathbf{x})]$ is defined in DPO where the normalizing constant term can be ignored as it will be canceled out in the DPO objective, which only involves the difference of the rewards for the same prompt. In this work, we explore whether the above readily available implicit reward model after DPO training provides an opportunity to further improve the language model.
@@
This paper answers the research question in the affirmative, by using the above implicit rewards in a bootstrapping fashion to further improve the LLM alignment with human preferences. Specifically, our approach follows the iterative DPO framework [25], where the implicit rewards are provided as the preference signals, as illustrated in Figure 1. We start with a model that has been through one round of DPO using human preference data. We then use the implicit rewards induced by itself to rank outputs from the current LLM, thereby, yielding a new preference dataset cheaply. We run DPO again with this newly generated preference dataset to obtain an updated LLM and then repeat the process. However, the above approach still needs further refinement to address practical issues. One is the known issue of length exploitation [15] where LLMs generate long responses when the same content could be provided more succinctly. Another issue is that the implicit reward model is an approximate proxy for human preferences, hence relying on it strongly can result in corruption of the initial knowledge inbuilt into the LLM.
@@
Figure 1: (Left) Iterative DPO [25] with various preference signals: under the iterative DPO framework, the policy model is iteratively trained on a newly generated preference dataset. This dataset can be constructed using various preference signals. A common source is a scalar reward model (RM) [14], denoted as $r_{\phi}$. Alternatively, the dataset can be created by prompting a LLM to judge the responses. This LLM can either be an external model $\pi_{\phi}$ or the policy model itself from the previous iteration $\pi_{\theta^{(t-1)}}$. In this context, ${\mathbf{x}}$ and ${\mathbf{y}}$ are processed through a LLM-as-a-judge template $g(\cdot,\cdot)$ [28]. We propose utilizing the length-regularized (LR) implicit rewards introduced in Section 4.2, where $\pi_{\theta^{(t-1)}}$ and $\pi_{\mathrm{ref}}^{(t-1)}$ represent the policy model and reference model from the previous DPO iteration, respectively. Our experiment excludes approaches requiring external models, such as $r_{\phi}$ and $\pi_{\phi}$, as they are beyond this work’s scope. (Right) Our method, which leverages implicit rewards, further improves DPO-tuned models by a large margin, resulting in superior performance compared to the prompting counterpart.
@@
We address the length exploitation issue by length-regularized reward shaping, which discourages long responses from being preferred. To fix the overreliance on the implicit reward we use insights from continual learning [18] and replay high quality human preference data that was used in the first round of DPO (this round was before we start our bootstrapping). Our method, named self-alignment with DPO ImpliCit rEwards (DICE), significantly improves LLM alignment quality with different base models. On AlpacaEval 2, we achieve $8.02\%$ length-controlled win rate improvement with the Zephyr-based model and $9.35\%$ improvement with the Llama3-based model. Our best model outperforms Gemini Pro, but with only 8B parameters and does not require any in-house data or external reward model.
@@
  • We propose to utilize the implicit reward model readily available in a DPO-ed LLM. The implicit reward model enables us to evaluate the responses generated by the current policy and construct a preference dataset for future rounds of DPO without any external feedback;
@@
  • Empirical results show that our approach DICE enables significant improvement in alignment with different base models, and our best bootstrapped model achieves better performance than Gemini Pro but with only 8B parameters.
2 Related Work
Self-Improving Fine-Tuning. Many efforts have been made to investigate ways of fine-tuning language models without a large amount of human annotation [9, 11, 20, 21, 28]. Starting from an SFT model, Sun et al. [20] collect the preference labels by prompting the SFT model itself to choose the preferred response given a principle and train a principle-driven reward model with these preference labels. Afterwards, they optimize the policy by PPO with their reward model. Yuan et al. [28] construct the preference dataset by their own supervised fine-tuned model trained on instruction following data and evaluation fine-tuning data, followed by DPO training on the preference dataset. Our work differs from this line of work in the motivations and assumptions. This line of work aims to align a language model with a small amount of seed data by eliciting the internal knowledge that is learned during the pretraining phase of LLMs. In contrast, our goal is to further improve a DPO-ed model in a bootstrapping manner by utilizing its implicit rewards.
On-Policy Sampling in Preference Tuning. DPO and its variants are popular due to their simplicity in training and implementation. However, research indicates that the offline nature of these direct alignment from preference (DAP) algorithms often prevents them from learning a good policy [8]. Guo et al. [8] have shown that the offline DPO will quickly overfit the preference dataset, while it performs much better and is more stable if online feedback can be provided to their on-policy samples [^1] On-policy samples: the data collected while following the current policy that is being optimized.. Tajwar et al. [23] systematically discussed the properties of different preference fine-tuning approaches and the role of on-policy sampling. Their actionable takeaways for practitioners tell on-policy sampling generally improves performance and efficiency. If not being able to perform the pure on-policy sampling, using the data that is closer to on-policy samples also helps as it can be seen in iterative DPO [25] and self-rewarding language models [28]. Similarly, our approach enables us to train the policy model with the preference data which is closer to on-policy samples than the offline dataset without any external reward models. We hypothesize this is one of the main gain sources of our approach.
DPO Implicit Rewards. Rafailov et al. [16] recently study DPO from a token-level MDP perspective, and reveal that DPO-ed models implicitly parameterize token-wise dense reward functions. They therefore conduct beam search using the implicit rewards to improve the inference quality. Our work is inspired by this observation but focuses on utilizing the DPO-ed model’s implicit rewards to bootstrap itself for self-alignment. Another independent work to ours [30] pretrains a DPO model to serve as a standalone dense reward generator for PPO training, instead of further improving the DPO model.
@@ 3 Preliminaries
We provide a brief review of the standard RLHF [14] and DPO algorithms [17]. Through the review, we demonstrate the implicit reward model induced by DPO, which will be used in our work.
In preference tuning, the preference data typically takes the form of pairwise preferences. Each prompt ${\mathbf{x}}$ is paired with two possible responses, ${\mathbf{y}}_{1}$ and ${\mathbf{y}}_{2}$. The human annotator [3] or AI annotator [10] provides the preference feedback $o(\mathbf{y}_{1}\succ\mathbf{y}_{2}|\mathbf{x})\in\{0,1\}$, indicating whether $\mathbf{y}_{1}$ is preferred over $\mathbf{y}_{2}$. The preferred response is denoted as ${\mathbf{y}}_{w}$, while the other is denoted as ${\mathbf{y}}_{l}$. A common assumption is that the ground-truth human preferences follow the Bradley-Terry model [2]. Based on this assumption, we can train a parameterized reward model $r_{\phi}(\mathbf{x},\mathbf{y})$ using maximum likelihood:
@@ 3.1 Reinforcement Learning from Human Feedback
where $\pi_{\textrm{ref}}(\mathbf{y}|\mathbf{x})$ denotes a reference distribution, and $\beta$ is a hyper-parameter. This objective balances the maximization of the reward $r_{\phi}(\mathbf{x},\mathbf{y})$ and divergence from the fixed reference distribution. The divergence term, given by the KL divergence (i.e., $\mathbb{D}_{\mathrm{KL}}\left[\pi_{\theta}(\mathbf{y}|\mathbf{x})\|\pi_{\textrm{ref}}(\mathbf{y}|\mathbf{x})\right]$) acts as a regularizer to prevent the policy $\pi_{\theta}$ from drifting too far away from the initial distribution $\pi_{\textrm{ref}}(\mathbf{y}|\mathbf{x})$. This objective is then optimized using a general-purpose RL algorithm, such as PPO [19].
@@ 3.2 Direct Preference Optimization
DPO [17] starts with the same objective as Eq. 2. However, DPO derives an analytical closed-form relation between the reward and the resulting optimal policy. This relation can be used to reparameterize the ground truth reward in terms of the corresponding optimal policy. This reparameterized formulation can be substituted back into the reward optimization objective in Eq. 1, enabling direct training of the optimal model on the feedback data using the following objective:
@@
This reward function is commonly referred to as an “implicit reward” [16, 30]. Theorem 1 in Rafailov et al. [17] demonstrates that this parameterization of a reward model is indeed valid without loss of generality. If we substitute this form of $r_{\theta}(\mathbf{x},\mathbf{y})$ into the RL objective in Eq. 2, we can derive the optimal solution in a closed form, which is $\pi_{\theta}$. Consequently, once DPO optimization is completed, we obtain an “implicit reward model” as defined by Eq. 4.
@@ 4 Bootstrapping with DPO Implicit Rewards
DPO is an attractive alternative to RLHF as it largely simplifies the implementation and training process of language model alignment. However, recent evidences [8, 25] have shown that continuing DPO training on a fixed offline dataset results in inferior policy, while the policy can be further improved if one can collect new responses generated by the updated policy and provide preference labels to perform another round of DPO training. This can be understood as being closer to the on-policy sampling, which is generally preferred in preference fine-tuning [23, 24]. In Section 4.1, we first outline the iterative DPO training framework adopted in this work, and provide a theoretical analysis similar to Xie et al. [27] to show that learning with on-policy samples can be more effective than utilizing an offline dataset. Next, in Section 4.2, we introduce the proposed Length-Regularized Implicit Rewards, which augment the vanilla implicit rewards with a length-regularized reward shaping, to judge the on-policy sampled responses for constructing the preference dataset. Notably, by fine-tuning a DPO-ed LM on the constructed dataset, we essentially align it without relying on any external preference feedback (e.g., RLHF or RLAIF), hence in a bootstrapping fashion. Furthermore, to mitigate the potential catastrophic forgetting in the continual fine-tuning, we propose experience replay (Section 4.3) that mixes the generated data with the offline data for better performance. We refer to our method as iterative self-alignment (bootstrapping) with DPO ImpliCit rEwards (DICE).
4.1 Iterative DPO with On-policy Sampling
@@
We employ the iterative DPO preference tuning framework, where we start from a base language model (a base policy) $\pi_{\theta^{(0)}}$ that is DPO-tuned from an initial reference model $\pi_{\text{ref}}^{(0)}$, commonly an SFT model. In each round $t\in\{1,2,\dots\}$, we sample $K$ on-policy responses from the latest policy $\pi_{\theta^{(t-1)}}(\cdot\mid{\mathbf{x}})$ given a prompt ${\mathbf{x}}$. We then label the response with the highest and the lowest implicit rewards as winning and losing responses respectively, thus constructing a new preference dataset ${\mathcal{D}}_{t}$. We further fine-tune the policy with DPO’s objective (Eq. 3) to obtain the updated policy $\pi_{\theta^{(t)}}$ with reference model $\pi^{(t)}_{\text{ref}}=\pi_{\theta^{(t-1)}}$. This process is repeated to iteratively improve the language model.
To investigate the effect of on-policy sampling, we make a generalized notation for the sampling policy at the $t$-th round as $\pi^{(t)}$, and compare the on-policy sampling ($\pi^{(t)}=\pi_{\theta^{(t-1)}}$) with sampling from an offline dataset ($\pi^{(t)}=\pi_{\mu}$). For a prompt ${\mathbf{x}}$, we denote its optimal response as ${\mathbf{y}}^{\star}$ and suboptimal ones as ${\mathcal{S}}=\{{\mathbf{y}}^{-}_{i}\}$. We also denote its preferences sampled from $\pi^{(t)}$ as $({\mathbf{y}}^{(t)}_{w},{\mathbf{y}}^{(t)}_{l})$. By Eq. 3 and the definition of the logistic function, we can rewrite the DPO loss at round $t$ for the sample $({\mathbf{x}},{\mathbf{y}}^{(t)}_{w},{\mathbf{y}}^{(t)}_{l})$ as:
@@
Figure 2: Distribution of the length difference between the winning and losing examples ($|\mathbf{y}_{w}|-|\mathbf{y}_{l}|$). (Top) Distribution of the first round on-policy generated dataset. With LR reward shaping defined by Eq. 8, the length bias is mitigated and the length difference becomes more evenly distributed. The average length difference decreases from $1031$ to $-21$ by setting $\alpha$=$0.023$. (Bottom) Distribution of the high quality UltraFeedback preference dataset [4] is almost unbiased.
@@ 4.2 Length-Regularized Implicit Rewards
It is a known issue in the literature that preference tuning may introduce length bias (or length exploitation) [15] , which is likely caused by the fact that the preference labels collected from human annotators favor more verbose responses. This problem is further compounded by the iterative self-alignment scheme such as the one in Yuan et al. [28], because the generated responses that are long and preferred will be reinforced in the next round of DPO, leading the language model to generate increasingly longer responses.
Inevitably, the vanilla DPO implicit rewards as in Eq. 4 would also exhibit length bias when generating preference dataset. In Figure 2, we show the distribution of the difference in string length of the winning and losing responses. We can see from the top figure that vanilla implicit rewards yield a skewed distribution (in green), with an average length difference $1031$. In stark contrast, the length difference of a high quality preference dataset is almost normally distributed (as in the bottom figure). This observation motivates us to debias the distribution induced by vanilla implicit rewards so as to mitigate the length exploitation. We resort to reward shaping [22] for this purpose. In particular, we introduce a length-regularized (LR) reward shaping term in the implicit reward that penalizes the length of the response to obtain the shaped reward:
@@
$\alpha^{\star}=\operatorname*{arg\,min}_{\alpha}\mathbb{E}_{({\mathbf{y}}_{w},{\mathbf{y}}_{l})\sim{\mathcal{D}}(\alpha)}\left(\Big{|}|{\mathbf{y}}_{w}|-|{\mathbf{y}}_{l}|\Big{|}\right).$ (9)
We can employ any black-box optimizer to solve this non-differentiable objective function. In this work we find a simple random search suffices, and its solution effectively transforms the dataset into a more evenly distributed one (shown in the orange curve in the top of Figure 2). We will output ${\mathcal{D}}(\alpha^{\star})$ for the next round of DPO training. Details of the optimization can be found in Appendix B.
Importantly, despite the resemblance of Eq. 8 to Park et al. [15] where they incorporate the token length as a regularizer in the training objective, our reward shaping is conducted during the dataset construction stage, thereby avoiding the need for expensive hyper-parameter tuning.
@@
1:  Input: prompt set $\mathcal{X}$ extracted from ${\mathcal{D}}_{\mathrm{offline}}$, initial DPO-tuned policy $\pi_{{\theta}^{(0)}}$, initial reference policy $\pi^{(0)}_{\mathrm{ref}}$, number of generated samples $K$, regularization weight $\beta$, experience replay weight $\gamma\in(0,1)$
@@
4:     Create preference dataset ${\mathcal{D}}(\alpha^{\star})$ by optimizing Eq. 9;// ${\mathcal{D}}(\alpha)$ is constructed based on LR rewards $r_{\mathrm{LR}}(\mathbf{x},\mathbf{y}_{k};\alpha)$, $k\in[K]$;// Evaluate $r_{\mathrm{LR}}(\mathbf{x},\mathbf{y}_{k};\alpha)$ given $\pi_{{\theta}^{(t-1)}}$ and $\pi^{(t-1)}_{\mathrm{ref}}$ as the target policy and reference policy
@@
${\theta}^{(t)}\xleftarrow{}\operatorname*{arg\,min}_{\theta}-\mathbb{E}_{\left(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l}\right)\sim\mathcal{D}_{t}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}\left(\mathbf{y}_{w}|\mathbf{x}\right)}{\pi_{{\theta}^{(t-1)}}\left(\mathbf{y}_{w}|\mathbf{x}\right)}-\beta\log\frac{\pi_{\theta}\left(\mathbf{y}_{l}|\mathbf{x}\right)}{\pi_{{\theta}^{(t-1)}}\left(\mathbf{y}_{l}|\mathbf{x}\right)}\right)\right];$
7:     Assign $\pi^{(t)}_{\mathrm{ref}}\xleftarrow{}\pi_{{\theta}^{(t-1)}}$;
8:  end for

### v2

@@ Abstract
[…]
@@ 1 Introduction
Direct preference optimization (DPO) (Rafailov et al., 2024b) presents a compelling alternative to reinforcement learning from human feedback (RLHF) (Stiennon et al., 2020) in large language models (LLMs). By circumventing the complexity of learning a reward model from given human preferences, DPO is simpler to implement and train compared to the RLHF approaches. Importantly, DPO, once trained, implicitly specifies a reward model. Mathematically, the reward for a response $\mathbf{y}$ to the prompt $\mathbf{x}$ can be expressed in terms of the optimal policy $\pi^{\star}$ and the reference policy $\pi_{\textrm{ref}}$:
[…]
@@
for parameter $\beta$ and normalizing constant $Z$. Further, an implicit reward $r(\mathbf{x},\mathbf{y})=\beta\cdot[\log\pi^{\star}(\mathbf{y}|\mathbf{x})-$ $\log\pi_{\textrm{ref}}(\mathbf{y}|\mathbf{x})]$ is defined in DPO where the normalizing constant term can be ignored as it will be canceled out in the DPO objective, which only involves the difference of the rewards for the same prompt. In this work, we explore whether the above readily available implicit reward model after DPO training provides an opportunity to further improve the language model.
@@
This paper answers the research question in the affirmative, by using the above implicit rewards in a bootstrapping fashion to further improve the LLM alignment with human preferences. Specifically, our approach follows the iterative DPO framework (Tran et al., 2023), where implicit rewards serve as the preference signals, as illustrated in Figure 1. We start with a model that has been through one round of DPO using human preference data, referred to as a DPO-tuned model. We then use the implicit rewards induced by DPO itself to rank outputs from the current LLM, thereby, yielding a new preference dataset cheaply. We run DPO again with this newly generated preference dataset to obtain an updated LLM and then repeat the process. However, the above approach still needs further refinement to address practical issues. One is the known issue of length exploitation (Park et al., 2024) where LLMs generate long responses when the same content could be provided more succinctly. Another issue is that the implicit reward model is an approximate proxy for human preferences, hence relying on it strongly can result in corruption of the initial knowledge inbuilt into the LLM.
@@
Figure 1: (Left) Iterative DPO (Tran et al., 2023) with various preference signals: under the iterative DPO framework, the policy model is iteratively trained on a newly generated preference dataset. This dataset can be constructed using various preference signals. A common source is a scalar reward model (RM) (Ouyang et al., 2022), denoted as $r_{\phi}$. Alternatively, the dataset can be created by prompting a LLM to judge the responses. This LLM can either be an external model $\pi_{\phi}$ or the policy model itself from the previous iteration $\pi_{\theta^{(t-1)}}$. In this context, ${\mathbf{x}}$ and ${\mathbf{y}}$ are processed through a LLM-as-a-judge template $g(\cdot,\cdot)$ (Yuan et al., 2024). We propose utilizing the length-regularized (LR) implicit rewards introduced in Section 3.1, where $\pi_{\theta^{(t-1)}}$ and $\pi_{\mathrm{ref}}^{(t-1)}$ represent the policy model and reference model from the previous DPO iteration, respectively. Our experiment excludes approaches requiring external models, such as $r_{\phi}$ and $\pi_{\phi}$, as they are beyond this work’s scope. (Right) Our method, which leverages implicit rewards, further improves DPO-tuned models by a large margin, resulting in superior performance compared to the prompting counterpart.
@@
Inspired by (Park et al., 2024), we address the length exploitation issue by length-regularized reward shaping, which discourages long responses from being preferred. Notably different from Park et al. (2024), who incorporate response length into the loss function, we directly search for a length-unbiased dataset, avoiding the expensive hyper-parameter tuning. To fix the overreliance on implicit rewards, we leverage insights from continual learning (Rolnick et al., 2019) and replay high quality human preference data that was used in the first round of DPO (the round before bootstrapping began). Our method, named self-alignment with DPO ImpliCit rEwards (DICE), significantly improves LLM alignment quality with different base models. On AlpacaEval 2, we achieve $8.02\%$ length-controlled (LC) win rate improvement with the Zephyr-based model and $9.35\%$ improvement with the Llama3-based model.
@@
  • We propose to utilize the implicit reward model readily available in a DPO-tuned LLM. The implicit reward model enables us to evaluate the responses generated by the current policy and construct a preference dataset for future rounds of DPO without any external feedback;
@@
[…]
@@ 2 Preliminaries
We provide a brief review of the standard RLHF (Stiennon et al., 2020; Ouyang et al., 2022) and DPO algorithms (Rafailov et al., 2024b). Through the review, we demonstrate the implicit reward model induced by DPO, which will be used in our work.
In preference tuning, the preference data typically takes the form of pairwise preferences. Each prompt ${\mathbf{x}}$ is paired with two possible responses, ${\mathbf{y}}_{1}$ and ${\mathbf{y}}_{2}$. The human annotator (Christiano et al., 2017) or AI annotator (Lee et al., 2023) provides the preference feedback $o(\mathbf{y}_{1}\succ\mathbf{y}_{2}|\mathbf{x})\in\{0,1\}$, indicating whether $\mathbf{y}_{1}$ is preferred over $\mathbf{y}_{2}$. The preferred response is denoted as ${\mathbf{y}}_{w}$, while the other is denoted as ${\mathbf{y}}_{l}$. A common assumption is that the ground-truth human preferences follow the Bradley-Terry model (Bradley & Terry, 1952). Based on this assumption, we can train a parameterized reward model $r_{\phi}(\mathbf{x},\mathbf{y})$ using maximum likelihood:
@@ 2.1 Reinforcement learning from human feedback
where $\pi_{\textrm{ref}}(\mathbf{y}|\mathbf{x})$ denotes a reference distribution, and $\beta$ is a hyper-parameter. This objective balances the maximization of the reward $r_{\phi}(\mathbf{x},\mathbf{y})$ and divergence from the fixed reference distribution. The divergence term, given by the KL divergence (i.e., $\mathbb{D}_{\mathrm{KL}}\left[\pi_{\theta}(\mathbf{y}|\mathbf{x})\|\pi_{\textrm{ref}}(\mathbf{y}|\mathbf{x})\right]$) acts as a regularizer to prevent the policy $\pi_{\theta}$ from drifting too far away from the initial distribution $\pi_{\textrm{ref}}(\mathbf{y}|\mathbf{x})$. This objective is then optimized using a general-purpose RL algorithm, such as PPO (Schulman et al., 2017).
@@ 2.2 Direct preference optimization
DPO (Rafailov et al., 2024b) starts with the same objective as Eq. 2 but derives an analytical closed-form relation between the reward and the resulting optimal policy. This relation can be used to reparameterize the ground truth reward in terms of the corresponding optimal policy. This reparameterized formulation can be substituted back into the reward optimization objective in Eq. 1, enabling direct training of the optimal model on the feedback data using the following objective:
@@
This reward function is commonly referred to as an “implicit reward” (Rafailov et al., 2024a; Zhong et al., 2024). Theorem 1 in Rafailov et al. (2024b) demonstrates that this parameterization of a reward model is indeed valid without loss of generality. If we substitute this form of $r_{\theta}(\mathbf{x},\mathbf{y})$ into the RL objective in Eq. 2, we can derive the optimal solution in a closed form, which is $\pi_{\theta}$. Consequently, once DPO optimization is completed, we obtain an “implicit reward model” as defined by Eq. 4.
@@ 3 Bootstrapping with DPO implicit rewards
DPO is an attractive alternative to RLHF as it largely simplifies the implementation and training process of language model alignment. However, recent evidences (Guo et al., 2024; Tran et al., 2023) have shown that continuing DPO training on a fixed offline dataset results in inferior policy, while the policy can be further improved if one can collect new responses generated by the updated policy and provide preference labels to perform another round of DPO training. This can be understood as being closer to the on-policy sampling, which is generally preferred in preference fine-tuning (Tajwar et al., 2024; Tang et al., 2024).
@@
In this work, we employ such iterative DPO preference tuning framework, where we start from a base language model (a base policy) $\pi_{\theta^{(0)}}$ that is DPO-tuned from an initial reference model $\pi_{\text{ref}}^{(0)}$, commonly an SFT model. In each round $t\in\{1,2,\dots\}$, we sample $K$ on-policy responses from the latest policy $\pi_{\theta^{(t-1)}}(\cdot\mid{\mathbf{x}})$ given a prompt ${\mathbf{x}}$. We then label the response with the highest and the lowest implicit rewards as winning and losing responses respectively, thus constructing a new preference dataset ${\mathcal{D}}_{t}$. We further fine-tune the policy with DPO’s objective (Eq. 3) to obtain the updated policy $\pi_{\theta^{(t)}}$ with reference model $\pi^{(t)}_{\text{ref}}=\pi_{\theta^{(t-1)}}$. This process is repeated to iteratively improve the language model. Notably, by fine-tuning a DPO-tuned LM on the its own constructed dataset, we essentially align it without relying on any external preference feedback (e.g., RLHF or RLAIF), hence in a bootstrapping fashion. We thus refer to our method as iterative self-alignment (bootstrapping) with DPO ImpliCit rEwards (DICE).
We next introduce two ingredients that are proven critical in the self-alignment process. In Section 3.1, we present Length-Regularized Implicit Rewards, which augment the vanilla implicit rewards with a length-regularized reward shaping, to judge the on-policy sampled responses for constructing the preference dataset. Furthermore, to mitigate the potential catastrophic forgetting in the continual fine-tuning, we propose experience replay (Section 3.2) that mixes the generated data with the offline data for better performance.
@@
Figure 2: Distribution of the length difference between the winning and losing examples ($|\mathbf{y}_{w}|-|\mathbf{y}_{l}|$). (Top) Distribution of the first round on-policy generated dataset. With LR reward shaping defined by Eq. 5, the length bias is mitigated and the length difference becomes more evenly distributed. The average length difference decreases from $1031$ to $-21$ by setting $\alpha$=$0.023$. (Bottom) Distribution of the high quality UltraFeedback preference dataset (Cui et al., 2023) is almost unbiased.
@@ 3.1 Length-regularized implicit rewards
It is a known issue in the literature that preference tuning may introduce length bias (or length exploitation) (Park et al., 2024) , which is likely caused by the fact that the preference labels collected from human annotators favor more verbose responses. This problem is further compounded by the iterative self-alignment scheme such as the one in Yuan et al. (2024), because the generated responses that are long and preferred will be reinforced in the next round of DPO, leading the language model to generate increasingly longer responses.
Inevitably, the vanilla DPO implicit rewards as in Eq. 4 would also exhibit length bias when generating preference dataset. In Figure 2, we show the distribution of the difference in string length of the winning and losing responses. We can see from the top figure that vanilla implicit rewards yield a skewed distribution (in green), with an average length difference $1031$. In stark contrast, the length difference of a high quality preference dataset is almost normally distributed (as in the bottom figure). This observation motivates us to debias the distribution induced by vanilla implicit rewards so as to mitigate the length exploitation. We resort to reward shaping (Sutton & Barto, 2018) for this purpose. In particular, we introduce a length-regularized (LR) reward shaping term in the implicit reward that penalizes the length of the response to obtain the shaped reward:
@@
$\alpha^{\star}=\operatorname*{arg\,min}_{\alpha}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Big{|}}\mathbb{E}_{({\mathbf{y}}_{w},{\mathbf{y}}_{l})\sim{\mathcal{D}}(\alpha)}\left(|{\mathbf{y}}_{w}|-|{\mathbf{y}}_{l}|\right){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Big{|}}.$ (6)
We can employ any black-box optimizer to solve this non-differentiable objective function. In this work we find a simple random search suffices, and its solution effectively transforms the dataset into a more evenly distributed one (shown in the orange curve in the top of Figure 2). We will output ${\mathcal{D}}(\alpha^{\star})$ for the next round of DPO training. Details of the optimization can be found in footnote 10.
Importantly, despite the resemblance of Eq. 5 to Park et al. (2024) where they incorporate the token length as a regularizer in the training objective, our reward shaping is conducted during the dataset construction stage, thereby avoiding the need for expensive hyper-parameter tuning.
@@
1:  Input: prompt set $\mathcal{X}$ extracted from ${\mathcal{D}}_{\mathrm{offline}}$, initial DPO-tuned policy $\pi_{{\theta}^{(0)}}$, initial reference policy $\pi_{\mathrm{ref}}=\pi_{{\theta}^{(-1)}}$, number of generated samples $K$, regularization weight $\beta$, experience replay weight $\gamma\in(0,1)$
@@
4:     Create preference dataset ${\mathcal{D}}(\alpha^{\star})$ by optimizing Eq. 6;// ${\mathcal{D}}(\alpha)$ is constructed by taking the best and worst ${\mathbf{y}}$ based on $r_{\mathrm{LR}}(\mathbf{x},\mathbf{y}_{k};\alpha)$, $k\in[K]$;// Evaluate $r_{\mathrm{LR}}(\mathbf{x},\mathbf{y}_{k};\alpha)$ given $\pi_{{\theta}^{(t-1)}}$ and $\pi_{{\theta}^{(t-2)}}$ as the target policy and reference policy
@@
${\theta}^{(t)}\xleftarrow{}\operatorname*{arg\,min}_{\theta}-\mathbb{E}_{\left(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l}\right)\sim\mathcal{D}_{t}}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}\left(\mathbf{y}_{w}|\mathbf{x}\right)}{\pi_{{\theta}^{(t-1)}}\left(\mathbf{y}_{w}|\mathbf{x}\right)}-\beta\log\frac{\pi_{\theta}\left(\mathbf{y}_{l}|\mathbf{x}\right)}{\pi_{{\theta}^{(t-1)}}\left(\mathbf{y}_{l}|\mathbf{x}\right)}\right)\right]\text{;}$
7:  end for